OpenStudy (anonymous):

Express as a polynomial \[(a+b+c+d)^2\]

OpenStudy (zepp):

Use the algebraic identity \((a+b)^2=a^2+2ab+c^2\) \(((a+b)+(c+d))^2\) \((a+b)^2+2(a+b)(c+d)+(c+d)^2\)

OpenStudy (zepp):

Use the identity again, FOIL out the middle term; \(a^2+2ab+b^2+c^2+2cd+d^2+2(ac+ad+bc+bd)\)

OpenStudy (zepp):

\(a^2+b^2+c^2+d^2+2ab+2cd+2ac+2ad+2bc+2bd\)

OpenStudy (anonymous):

o.O... I'm trying to look beyond all the terms...

OpenStudy (zepp):

What do you mean? D:

OpenStudy (anonymous):

Whew, okay I got thank same thing thanks zepp!

OpenStudy (anonymous):

It just looks confusing... and hurts my eyes, lol...

OpenStudy (lgbasallote):

there's a shortcut to this if you're interested :) step 1: square first term = a^2 step 2: square second term = b^2 step 3: square third term = c^2 step 4: square fourth term = d^2 step 5: twice the product of first term and second term = 2ab step 6: twice the product of first term and third term = 2ac step 6: twice the product of first term and fourth term = 2ad step 7: twice the product of second term and third term = 2bc step 8: twice the product of second term and fourth term = 2bd step 9: twice the product of third and fourth term = 2cd step 10: combine everything = a^2 + b^2 + c^2 + d^2 + 2ab + 2ac + 2ad + 2bc + 2bd + 2cd this works for all square of polynomials

OpenStudy (lgbasallote):

here is a tutorial where i proved this http://openstudy.com/updates/4f9dda4ae4b000ae9ed25338 :)

OpenStudy (anonymous):

I am interested by so many steps...

OpenStudy (lgbasallote):

it's just all squares and twice of products lol

OpenStudy (anonymous):

I don't understand why you do step 5 to 9...

OpenStudy (anonymous):

in such manner I guess...

OpenStudy (anonymous):

OMG!!!... I see it, haha! thanks... that is beautiful @lgbasallote

OpenStudy (lgbasallote):

haha glad to hear :)